# OV-MER: Towards Open-Vocabulary Multimodal Emotion Recognition

### v2

@@ 4.1 CLUE Generation
To reveal the modality impact, we propose three variants of CLUE-Multi: CLUE-Audio, CLUE-Text, and CLUE-Video. In Figure 3, we illustrate their generation process. (1) CLUE-Audio: We observe that ALLM cannot fully leverage the text, and using an additional LLM to emphasize the text can further improve performance, which is also verified in Section 5. Therefore, we merge the checked acoustic clues with text using an additional LLM; (2) CLUE-Text: We only use the text to infer emotional states; (3) CLUE-Video: Since the visual content does not contain audio and text, we only use the checked visual clues. See Appendix P for more examples.
@@
MLLMs can address various multimodal tasks. Since emotion recognition relies on temporal information, we choose models that support at least video or audio. To generate CLUE-MLLM, we first use ALLM or VLLM to extract emotion-related descriptions, and then combine these descriptions with text using LLM. Compared with CLUE-Multi, this process does not use manually checked clues. Appendix Q provides model cards and relevant prompts. This paper aims to build a zero-shot benchmark for OV-MER, without the training process. All models are implemented in PyTorch, and all inference processes are carried out using a 32G NVIDIA Tesla V100 GPU.
@@
4.2 Metric Calculation
5 Results and Discussion
In this section, we default to using GPT-based grouping and employ GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as LLM. We generally report evaluation results in both languages, but if no specific language is mentioned, we default to reporting results for the English branch. To mitigate the impact of randomness, we conduct each experiment twice and report the average scores and standard deviations.
Main Results on CLUE-M/A/T/V.
For CLUE-M/A/T/V, most baselines use manually checked clues, serving as performance upper bounds of different modality combinations. In Table 2, we observe that CLUE-Multi performs the best, highlighting the importance of multimodal information in MER. In contrast, CLUE-Text performs the worst. This is because our OV-MERD dataset is derived from MER2023, where the contribution of text is smaller compared to audio and video [21]. Relying solely on text makes it difficult to recognize emotions accurately.
Main Results on CLUE-MLLM.
In Table 2, we introduce a heuristic baseline called Random, where we randomly select a label from basic emotions. This baseline reflects the lower bound of performance. We observe that MLLM generally outperforms Random, indicating that MLLM can partially address OV-MER. However, the performance of MLLM remains unsatisfactory, highlighting the limitations of existing MLLMs and the challenges of OV-MER. Furthermore, models that perform well in Chinese often perform well in English, suggesting that the impact of language differences on rankings is limited.
(a) Length (H)
(b) #Label (H)
(c) Word cloud (H)
(d) Length (H+L)
(e) #Label (H+L)
(f) Word cloud (H+L)
Figure 5: Human-only (H) vs. Human-LLM (H+L) strategy.
Model GPT M1 M2 M3-W1 M3-W2 M3-W3 M3-W4 M3-W5
Qwen-Audio 38.13±0.05 20.49±0.01 23.37±0.01 43.85±0.03 41.52±0.28 39.46±0.28 36.64±0.03 35.89±0.08
Otter 43.51±0.09 22.21±0.06 27.99±0.02 49.75±0.11 49.93±0.11 51.03±0.04 47.54±0.00 50.51±0.03
Video-LLaMA 44.73±0.14 23.56±0.08 28.39±0.17 52.90±0.12 53.60±0.04 47.50±0.20 52.97±0.09 46.39±0.12
VideoChat 45.53±0.11 22.15±0.00 26.24±0.08 47.79±0.07 47.76±0.11 46.78±0.11 49.53±0.15 45.93±0.18
SECap 45.72±0.09 26.52±0.01 32.88±0.03 52.26±0.03 52.11±0.03 50.77±0.03 50.43±0.03 49.97±0.03
Video-LLaVA 47.07±0.16 25.47±0.12 30.73±0.11 54.65±0.10 54.54±0.02 52.29±0.05 52.45±0.06 52.97±0.10
SALMONN 47.96±0.04 23.57±0.02 28.83±0.03 54.90±0.15 54.29±0.06 56.25±0.01 50.53±0.09 53.65±0.04
VideoChat2 49.07±0.26 26.92±0.09 31.40±0.10 52.38±0.13 53.56±0.13 52.14±0.23 50.63±0.19 51.37±0.14
Video-ChatGPT 50.52±0.06 28.99±0.04 34.05±0.05 57.66±0.04 57.37±0.00 55.50±0.13 55.24±0.02 52.93±0.05
LLaMA-VID 51.25±0.09 28.28±0.04 32.85±0.03 56.59±0.04 57.49±0.03 55.12±0.05 56.62±0.15 53.03±0.08
mPLUG-Owl 52.73±0.13 27.47±0.17 32.47±0.19 57.60±0.23 56.32±0.26 55.67±0.19 55.06±0.17 54.44±0.13
Chat-UniVi 53.08±0.01 28.89±0.02 33.23±0.08 57.00±0.06 57.50±0.03 56.80±0.01 55.86±0.07 55.81±0.02
PCC score 0.887 0.857 0.911 0.913 0.904 0.899 0.885
@@ 5 Results and Discussion
To verify the effectiveness of our human-LLM strategy, we additionally introduce a baseline using human-only annotation. In Figure 5, we compare two strategies from three aspects: the length distribution of generated descriptions, the distribution of sample-wise label numbers, and the word cloud. We observe that through human-LLM collaboration, we can obtain longer descriptions, provide more diverse labels for each sample, and generate a broader range of emotions. These results show that human-only annotation generally focuses on primary emotions while neglecting minor ones. With the pre-annotation and semantic reasoning capabilities of LLMs, we can obtain richer emotional labels. These results validate the effectiveness of our human-LLM collaborative strategy. Meanwhile, these results suggest that the LLM-driven approach does not lead to a narrow or biased interpretation of emotions, but rather helps uncover more subtle emotional nuances. We provide additional analysis in the Appendix R.
@@
We reveal the impact of different CLUE-MLLM generation strategies. Figure 6 introduces three methods: 1) S0 does not use text and inputs the video into MLLM; 2) S1 inputs both text and video into MLLM; 3) S2 first uses MLLM to extract descriptions and then combines with text using another LLM, same with the strategy in Figure 3. In Figure 4, S1 and S2 generally outperform S0, indicating the importance of the text content in OV-MER. Moreover, S2 typically performs better than S1. The reason is that inputting video and text into the MLLM simultaneously increases the task difficulty, and current MLLMs may struggle to handle complex prompts. S2 divides this process into two steps, reducing task complexity and achieving better performance. Therefore, we adopt S2 as the default strategy. More results are provided in Appendix Q.
GPT-based vs. EW-based Grouping.
We propose two grouping strategies: GPT-based and EW-based grouping. In this section, we explore the relationship between them. Table 3 reports $\mbox{F}_{\mbox{s}}$ for different EW-based strategies. For M3, we report the scores of the L1 strategy by default, and the complete results are provided in Appendix S. We observe that the PCC scores between GPT-based and EW-based groupings are relatively high, indicating that EW-based metrics can serve as an alternative to GPT-based metrics.
@@
In Table 2, CLUE-Multi performs the best. This leads to a hypothesis: Do sentences that are more “similar” to CLUE-Multi yield better emotion recognition performance? The most common way to measure “similarity” is through matching-based metrics, with BLEU1, BLEU4, METEOR, and ROUGE1 being the most widely used. Therefore, we use CLUE-MLLM as input and calculate both GPT-based and matching-based metrics (see Figure 7(a)), followed by calculating their PCC scores (see Figure 7(b)). From the experimental results, we have some interesting observations. First, the same metric across different languages typically exhibits high correlations. However, the correlation between GPT-based and matching-based metrics is not strong. For example, the highest PCC score between “F(E)” and matching-based metrics is 0.77. These results highlight the limitations of using matching-based metrics to evaluate the OV-MER task. The main reason is that matching-based metrics focus on low-level word-level matches, while emotion understanding is a relatively complex and high-level perceptual task. In Appendix T, we provide more examples for clarification.
@@
In this section, we conduct a user study to verify whether our dataset provides greater informativeness than basic emotions. Specifically, we hired four annotators and randomly selected samples from our dataset. For each sample, we provided both the basic emotion label and the OV-MERD label. The following instruction was given: Which label provides greater informativeness? The label with more information was marked as 1, and the other label as 0. Experimental results demonstrate that 97.5% of the results believed that our OV-MERD provided more information than basic emotions, verifying the effectiveness of our OV-MERD in emotion representation.
@@ 6 Conclusion
This paper extends traditional MER to OV-MER, allowing for the prediction of arbitrary numbers and types of emotions. To facilitate further research, we construct an initial dataset, define evaluation metrics, and propose solutions. We observe that current MLLMs struggle to achieve satisfactory results, as this task requires consideration of multimodal clues and subtle temporal changes, placing higher demands on MLLMs. Additionally, EW-based metrics can replace GPT-based metrics, thus reducing evaluation costs while ensuring reproducibility. This paper advances current research from basic to nuanced emotion recognition, which is crucial for building emotion AI.
Ethics Statement
The raw data of the OV-MERD dataset comes from the MER2023 dataset, from which we evenly select some samples with further annotation. Therefore, we do not collect new data; we just re-annotate existing data. This annotation process has received consent from the dataset owners and has passed our internal review. During the annotation process, we generously pay each annotator approximately ¥3,000 (around \$280), which is considered high. After proofreading our annotation results, we find that the annotations focus on the multimodal clues present in the videos, without any discriminatory annotations. Additionally, we restrict the use of the OV-MERD dataset to non-commercial purposes under the CC BY-NC 4.0 license. This license clearly outlines the correct and responsible use of our dataset. Details of our license are provided in the supplementary materials. In summary, this paper does not involve any ethical issues.
@@ Appendix A Task Comparison
One-Hot MER (OH-MER) is most widely discussed in the community. In this task, each sample is labeled with only one emotion. We should design a framework to predict the most likely label from a predefined emotional taxonomy [22]. Current research primarily focuses on the framework design and multimodal fusion strategy. For the former, research has shifted from traditional classifiers (e.g., SVM [23]) to deep models (e.g., Transformers [24]). For the latter, researchers mainly focus on how to align heterogeneous features (e.g., word alignment [25] and implicit alignment [26]) for subsequent multimodal fusion.
Multi-Label MER (ML-MER) considers the complexity of emotions (i.e., multiple emotions often occur simultaneously) and allows the model to predict multiple labels. The most straightforward solution is to use binary classifiers for each emotion [27, 28]. However, this approach is based on the assumption of emotion independence, ignoring the correlations between different emotions. Therefore, [29] and [30] proposed a sequence-to-emotion model, transforming the multi-label emotion recognition task into an emotion sequence prediction task, to further consider the correlations between emotions.
Partial Label Learning MER (PLL-MER) differs from the tasks mentioned above. In PLL-MER, each sample is associated with a set of candidate labels, only one of which is correct [31]. Research on PLL-MER can be roughly divided into two categories: average-based methods and identification-based approaches. The former assumes that each candidate label has an equal probability of being the ground truth [32], while the latter directly identifies the ground truth and maximizes its estimated probability [33].
Fine-Grained MER (FG-MER) differs from the above tasks in the label space. Previous tasks typically rely on coarse-grained emotion taxonomies, such as the widely accepted basic emotion taxonomies, like Ekman [6] or Plutchik [20] emotions. Differently, FG-MER aims to use fine-grained emotion taxonomies to capture more subtle emotions [34]. After expanding the label space, FG-MER follows the typical solution of OH-MER.
@@
  • Emotional Complexity. Human emotional states are diverse and nuanced. As psychologist Plutchik pointed out, humans can express approximately 34,000 different emotions [20]. Predefined taxonomies that categorize the full spectrum of emotions into a limited set of labels inevitably overlook some subtle emotional states. In contrast, OV-MER allows the model to understand and predict any emotion, enabling a more accurate modeling of the complex nature of human emotions.
  • Relationship Between OV-MER and ML-MER. Theoretically, ML-MER can be converted to OV-MER by spanning the label space to a complete set that includes all the emotion labels in the language (e.g., around 34,000 different emotions [20]). However, it is not feasible in practice to construct such a kind of dataset, i.e., annotators need to label each sample with 34,000 different emotions, and it’s hard to collect sufficient samples for every emotion category, let alone multiple labels).
@@
Appendix B Limitations
Firstly, the main contribution of this paper is the definition of a new task and the conduct of foundational research. In the future, we plan to design more effective frameworks to solve OV-MER. Specifically, we will incorporate more emotion-related instruction datasets to finetune MLLMs, thereby enhancing their emotion recognition ability. Meanwhile, how to integrate subtitle information and fuse multimodal inputs also plays a crucial role. We will also consider these aspects in the framework design.
Secondly, we evaluate some MLLMs, but not all models are covered. In the future, we will expand the scope of evaluation to cover more MLLMs to enrich our benchmark.
Thirdly, this paper does not involve cultural differences. Specifically, our original data is in Chinese, and the annotators we hired are also native Chinese speakers. In the future, we will also try to extend our method to other cultures and further analyze cultural differences.

### v3

@@ 4.1 CLUE Generation
To reveal the modality impact, we propose three variants of CLUE-Multi: CLUE-Audio, CLUE-Text, and CLUE-Video. In Figure 3, we illustrate their generation process. (1) CLUE-Audio: We observe that ALLM cannot fully leverage the text, and using an additional LLM to emphasize the text can further improve performance, which is also verified in Section 5. Therefore, we merge the checked acoustic clues with text using an additional LLM; (2) CLUE-Text: We only use the text to infer emotional states; (3) CLUE-Video: Since the visual content does not contain audio and text, we only use the checked visual clues. See Appendix Q for more examples.
@@
MLLMs can address various multimodal tasks. Since emotion recognition relies on temporal information, we choose models that support at least video or audio. To generate CLUE-MLLM, we first use ALLM or VLLM to extract emotion-related descriptions, and then combine these descriptions with text using LLM. Compared with CLUE-Multi, this process does not use manually checked clues. Appendix R provides model cards and relevant prompts. For MLLMs, we use their 7B version by default. All models are implemented in PyTorch, and all inference processes are executed on a 32GB NVIDIA Tesla V100 GPU.
4.2 Metric Calculation
5 Results and Discussion
In this section, we default to using GPT-based grouping and employ GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as LLM. We generally report evaluation results in both languages, but if no specific language is mentioned, we default to reporting results for the English branch. To mitigate the impact of randomness, we conduct each experiment twice and report the average scores and standard deviations. In addition to MLLM-based generative models, we also report the performance of discriminative models in Appendix U.
Main Results on CLUE-M/A/T/V.
For CLUE-M/A/T/V, most baselines use manually checked clues, serving as performance upper bounds of different modality combinations. In Table 2, we observe that CLUE-Multi performs the best, highlighting the importance of multimodal information in MER. Meanwhile, CLUE-Video outperforms CLUE-Text, consistent with the nature of our OV-MERD dataset. To be specific, OV-MERD is derived from MER2023, where the textual modality contributes less than the visual modality in emotion recognition (Lian et al., 2024b). Relying solely on text makes it difficult to recognize emotions accurately. Furthermore, CLUE-Audio achieves superior performance over both CLUE-Text and CLUE-Video, suggesting that although textual expressions may be ambiguous for emotion recognition, combining them with audio cues can effectively resolve these ambiguities.
Main Results on CLUE-MLLM.
In Table 2, we introduce a heuristic baseline called Random, where we randomly select a label from basic emotions. This baseline reflects the lower bound. We observe that MLLM generally outperforms Random, indicating that MLLM can partially address the OV-MER task. However, the performance of MLLM remains unsatisfactory, highlighting the limitations of existing MLLMs and the challenges of OV-MER. Furthermore, models that perform well in Chinese often perform well in English, suggesting that the impact of language differences on rankings is limited. Appendix S presents quantitative analysis results on language differences.
(a) Length (H)
(b) #Label (H)
(c) Word cloud (H)
(d) Length (H+L)
(e) #Label (H+L)
(f) Word cloud (H+L)
Figure 4: Human-only (H) vs. Human-LLM (H+L) strategy.
@@ 5 Results and Discussion
To verify the effectiveness of our human-LLM strategy, we additionally introduce a baseline using human-only annotation. In Figure 4, we compare two strategies from three aspects: the length distribution of generated descriptions, the distribution of sample-wise label numbers, and the word cloud. In Figure 4, we observe that through human-LLM collaboration, we can obtain longer descriptions, provide more diverse labels for each sample, and generate a broader range of emotions. These results demonstrate that human-only annotation generally focuses on primary emotions while neglecting minor ones. With the pre-annotation and semantic reasoning capabilities of LLMs, we can obtain richer emotional labels. These results validate the effectiveness of our human-LLM collaborative strategy. Meanwhile, these results suggest that the LLM-driven approach does not lead to a narrow or biased interpretation of emotions, but rather helps uncover more subtle emotional nuances. We provide additional analysis in Appendix T.
@@
We reveal the impact of different CLUE-MLLM generation strategies. Figure 6 introduces three methods: 1) S0 does not use text and inputs the video into MLLM; 2) S1 inputs both text and video into MLLM; 3) S2 first uses MLLM to extract descriptions and then combines with text using another LLM, same with the strategy in Figure 3. In Figure 5, S1 and S2 generally outperform S0, indicating the importance of the text content in OV-MER. Moreover, S2 typically performs better than S1. The reason is that inputting video and text into the MLLM simultaneously increases the task difficulty, and current MLLMs may struggle to handle complex prompts. S2 divides this process into two steps, reducing task complexity and achieving better performance. Therefore, we adopt S2 as the default strategy. More results are provided in Appendix R.
@@
In Table 2, CLUE-Multi demonstrates the best performance, leading to the hypothesis: Do sentences that are more similar to CLUE-Multi yield better emotion recognition performance? The most common way to measure “similarity” is through matching-based metrics, with BLEU1, BLEU4, METEOR, and ROUGE1 being the most widely used. To investigate this, we use CLUE-MLLM as input and calculate both GPT-based and matching-based metrics (see Figure 7(a)), and report their Pearson Correlation Coefficient (PCC) scores in Figure 7(b). Experimental results reveal several interesting observations. First, the same metric across different languages typically shows high correlations. However, the correlation between GPT-based and matching-based metrics is relatively weak. For instance, the highest PCC score between “F(E)” and matching-based metrics is only 0.77. This discrepancy arises because matching-based metrics focus on low-level word-level matches, whereas emotion understanding is a more complex, high-level perceptual task. Appendix V provides a more detailed explanation of these findings.
Model GPT M1 M2 M3-W1 M3-W2
L1 L2 L1 L2
PCC score 0.887 0.857 0.911 0.940 0.913 0.942
Model M3-W3 M3-W4 M3-W5 M-avg
L1 L2 L1 L2 L1 L2
PCC score 0.904 0.927 0.899 0.922 0.885 0.894 0.942
GPT-based vs. EW-based Grouping.
We propose two grouping strategies: GPT-based and EW-based grouping. In this section, we explore the relationship between them. Table 3 reports $\mbox{F}_{\mbox{s}}$ for different EW-based strategies, as this metric is used for the final ranking, and we compute the PCC scores between different metrics. We observe that the PCC scores between GPT-based and EW-based groupings are relatively high, indicating that EW-based metrics can serve as an alternative to GPT-based metrics. Meanwhile, we observe that M3-L2 is always more correlated with the GPT-based metrics than M3-L1. M3-L1 emphasizes coarse-grained clustering information, whereas M3-L2 emphasizes fine-grained clustering information. The higher correlation between M3-L2 and GPT-based metrics suggests that GPT-based metrics primarily rely on fine-grained emotion clustering during the metric calculation.
@@
We conducted a user study to evaluate whether our annotation manner provides greater informativeness compared to traditional basic emotions. Specifically, we recruited four annotators and randomly selected 20 samples from our dataset. For each sample, we presented both the basic emotion label and the OV-MERD label. Annotators were instructed as follows: Which label provides greater informativeness? The label with more information was marked as 1, and the other label as 0. Experimental results show that 97.5% of the annotations favored our OV-MERD labels, confirming their superiority in informativeness over basic emotions and verifying the effectiveness of our OV-MERD in emotion representation.
Alignment with Human Perception.
To evaluate how well OV-MER aligns with human perception, we conducted an additional user study. Specifically, we recruited nine annotators and randomly selected 20 samples from our dataset. Each annotator was presented with (sample, OV-MERD label) pairs and asked to judge their alignment with human perception using a binary (Yes/No) response format. To ensure annotation quality, we included inspection data consisting of (sample, incorrect label) pairs. The results show that 96% of the annotations confirmed the alignment between OV-MERD labels and human perception. Considering potential annotator errors, this result demonstrates that our OV-MERD labels align well with human perception.
@@ 6 Conclusion
This paper extends traditional MER to OV-MER, allowing for the prediction of arbitrary numbers and types of emotions. For this task, we build a dataset, define metrics, and propose solutions. We observe that current MLLMs struggle to achieve satisfactory results, as this task requires consideration of multimodal clues and subtle temporal changes, placing higher demands on MLLMs. Additionally, EW-based metrics can replace GPT-based metrics, thus reducing evaluation costs while ensuring reproducibility. This paper advances current research from basic to nuanced emotion recognition, which is crucial for emotion AI.
Acknowledgments
This work is supported by the Excellent Youth Program of State Key Laboratory of Multimodal Artificial Intelligence Systems (MAIS2024311), the National Natural Science Foundation of China (62201572, 62322120, 61831022, 62276259, U21B2010, 62271083, 62306316, 62176165, 62206136, 62476146), the Internal Research Project of Xi’an Jiaotong-Liverpool University (RDF-24-01-016), the Young Elite Scientists Sponsorship Program by CAST (2024QNRC001), and the University of Oulu& Research Council of Finland Profi 7 (352788).
Impact Statements
Ethics Statement.
The raw data of the OV-MERD dataset comes from MER2023, from which we select some samples with further annotation. Therefore, we do not collect new data; we just re-annotate existing data. This annotation process has received consent from the dataset owners and has passed our internal review. During the annotation process, we generously pay each annotator approximately ¥3,000 (around \$280), which is considered high. After proofreading our annotation results, we find that the annotations focus on the multimodal clues present in the videos, without any discriminatory annotations. Additionally, we restrict the use of the OV-MERD dataset to non-commercial purposes under the CC BY-NC 4.0 license. This license clearly outlines the correct and responsible use of our dataset.
Proper Use.
MER is a widely discussed research topic. In this paper, we extend traditional MER by providing more accurate emotion annotations that go beyond the fixed emotion taxonomy. In the license we provide, we restrict the use of this dataset to academic research; commercial usage is prohibited. Meanwhile, this dataset can only be used in non-sensitive human-computer interaction scenarios to enhance the machine’s ability to understand human emotions and respond appropriately. It cannot be used in sensitive areas. For example, in interrogation scenarios, emotion recognition results should not be used to determine whether a criminal has committed a crime; Emotion recognition results should also not be used in recruitment and loan approval processes, as this may lead to unfair treatment of certain groups and affect their employment and financial opportunities; In the field of education, emotion recognition results should not be used to evaluate the performance of teachers and students.
@@ Appendix A Task Comparison
One-Hot MER (OH-MER) is most widely discussed in the community. In this task, each sample is labeled with only one emotion. We should design a framework to predict the most likely label from a predefined emotional taxonomy (Zhang et al., 2024). Current research primarily focuses on the framework design and multimodal fusion strategy. For the former, research has shifted from traditional classifiers (e.g., SVM (Bishop, 2006)) to deep models (e.g., Transformers (Vaswani et al., 2017)). For the latter, researchers mainly focus on how to align heterogeneous features (e.g., word alignment (Gu et al., 2018) and implicit alignment (Tsai et al., 2019a)) for subsequent multimodal fusion.
Multi-Label MER (ML-MER) considers the complexity of emotions (i.e., multiple emotions often occur simultaneously) and allows the model to predict multiple labels. The most straightforward solution is to use binary classifiers for each emotion (Godbole & Sarawagi, 2004; He & Xia, 2018). However, this approach is based on the assumption of emotion independence, ignoring the correlations between different emotions. Therefore, Huang et al. (2021) and Deng & Ren (2020) proposed a sequence-to-emotion model, transforming the multi-label emotion recognition task into an emotion sequence prediction task, to further consider the correlations between emotions.
Partial Label Learning MER (PLL-MER) differs from the tasks mentioned above. In PLL-MER, each sample is associated with a set of candidate labels, only one of which is correct (Feng et al., 2020). Research on PLL-MER can be roughly divided into two categories: average-based methods and identification-based approaches. The former assumes that each candidate label has an equal probability of being the ground truth (Cour et al., 2009), while the latter directly identifies the ground truth and maximizes its estimated probability (Lv et al., 2020).
Fine-Grained MER (FG-MER) differs from the above tasks in the label space. Previous tasks typically rely on coarse-grained emotion taxonomies, such as the widely accepted basic emotion taxonomies, like Ekman (Ekman, 1992) or Plutchik (Plutchik, 1980) emotions. Differently, FG-MER aims to use fine-grained emotion taxonomies to capture more subtle emotions (Demszky et al., 2020). After expanding the label space, FG-MER follows the typical solution of OH-MER.
@@
  • Emotional Complexity. Human emotional states are diverse and nuanced. As psychologist Plutchik pointed out, humans can express approximately 34,000 different emotions (Plutchik, 1980). Predefined taxonomies that categorize the full spectrum of emotions into a limited set of labels inevitably overlook some subtle emotional states. In contrast, OV-MER allows the model to understand and predict any emotion, enabling a more accurate modeling of the complex nature of human emotions.
  • Relationship Between OV-MER and ML-MER. Theoretically, ML-MER can be converted to OV-MER by spanning the label space to a complete set that includes all the emotion labels in the language (e.g., around 34,000 different emotions (Plutchik, 1980)). However, it is not feasible in practice to construct such a kind of dataset, i.e., annotators need to label each sample with 34,000 different emotions, and it’s hard to collect sufficient samples for every emotion category, let alone multiple labels).
@@
Appendix B Limitations and Future Work
Firstly, the main contribution of this paper is the definition of a new task and the conduct of foundational research. In the future, we plan to design more effective frameworks to solve OV-MER. Specifically, we will incorporate more emotion-related instruction datasets to fine-tune MLLMs, thereby enhancing their emotion recognition ability. Meanwhile, how to integrate subtitle information and fuse multimodal inputs also plays a crucial role. We will also consider these aspects in the framework design. Secondly, we evaluate some MLLMs, but not all models are covered. In the future, we will expand the scope of evaluation to cover more MLLMs to enrich our benchmark. Thirdly, this paper does not involve cultural differences. Specifically, our original data is in Chinese, and the annotators we hired are also native Chinese speakers. In the future, we will also try to extend our method to other cultures and further analyze cultural differences. Fourth, we will focus on fairness-aware and unbiased modeling in future work. Fifth, OV-MERD is derived from MER2023, which is sourced from high-rated movies and TV shows. The high ratings serve as an implicit validation of the actors’ performances, ensuring spontaneous and realistic emotional expressions. Currently, this type of dataset is the mainstream in the MER research community, as it provides a cost-effective means to expand the dataset scale. In the future, we plan to apply for additional funding to collect data featuring spontaneous, real-life emotional expressions by recruiting participants. Furthermore, we will employ domain adaptation techniques (e.g., domain adversarial neural networks) to address potential domain gaps between different data sources.